Find the equation of the normal to the parabola y²+4x=0 at the point where the line y=x+c touches it.
1
Expert's answer
2019-07-12T12:20:06-0400

1. Find inverse functions of the line and the parabola:

"y = x + C \u21d2 x = y - C;"

"y^2+4x=0 ⇒x = -\frac{y^2}{4}."

 

2. The slope of the line with respect to x-axis:

"x = y - C = 1 * y - C;"

"k=1."

 

3. The tangent to the parabola and the line have the same slope, sо

"x = -\frac{y^2}{4};"

"x'(y) = -\frac{y}{2} = k = 1;"

"y = -2;"

"x = -\frac{y^2}{4} = -\frac{(-2)^2}{4} = -1."

"C=y-x=-1."

Thus, the line touches the parabola at the point (-1,-2).


4.The tangent and normal vectors of the line with the slope 'k' are defined by formulas respectively (in reverse order (y,x)):

"T'=(1,k)=(1,1);"

"N'=(-k,1)=(-1,1)."

In general order (x,y) we obtain:

"T=(1,1);"

"N=(1,-1)."


5. Hence, the equation of the normal is

"\frac{x + 1}{1} = \frac{y + 2}{-1};"

"x + 1 = -y - 2;"

"y = -x - 3."

"Answer: y = -x - 3."